OpenStudy (anonymous):

Factor the expression. c2 + 4c – 4

OpenStudy (michele_laino):

the requested factorization, is given by the subsequent formula: \[\left( {c - {c_1}} \right)\left( {c - {c_2}} \right)\] where c_1 and c_2 are the zeroes of the quadratic equation: \[{c^2} + 4c - 4 = 0\]

OpenStudy (anonymous):

@Boss_Avii this website should help https://mathway.com/

OpenStudy (johnweldon1993):

Well we can see it wont factor nicely So we can either use the quadratic formula, or complete the square